OpenStudy (anonymous):

What is the solution to the equation 7^–3x ≈ 9 ?

OpenStudy (anonymous):

\[-3x=\frac{ln(9)}{\ln(7)}\] is the first step

OpenStudy (anonymous):

oh I see so after doing ln(9)/ln(7) I'd set it equal to -3x and solve from there?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\frac{\ln(9)}{\ln(7)}\] is a number solve \[-3x=\frac{\ln(9)}{\ln(7)}\] by \[x=-\frac{\ln(9)}{3\ln(7)}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw